The Function N(x)= -0.0534x^2 + 0.337x + 0.969
gives the number of active-duty military personnel in the United States Army (in millions) for the years 1965-1972, where x=0 corresponds to 1965, x = 1 corresponds to 1966 and so on. For this period, when was the Army's personnel strength level at its highest, and what was it?
I put it into the equation x = -b/2a which I got -3.155 rounded off =3
i then put the -3 into the original equation and I got the answer of -.5226 so I am guessing I did something wrong!
Any help? Please and thank you! :)

Answer by stanbon(57387)

The Function N(x)= -0.0534x^2 + 0.337x + 0.969
gives the number of active-duty military personnel in the United States Army (in millions) for the years 1965-1972, where x=0 corresponds to 1965, x = 1 corresponds to 1966 and so on. For this period, when was the Army's personnel strength level at its highest, and what was it?
I put it into the equation x = -b/2a which I got -3.155 rounded off =3
i then put the -3 into the original equation and I got the answer of -.5226 so I am guessing I did something wrong!
----------------------------------------------------
For x = -b/2a you should get x = +3.155 which corresponds to the year
1965 + 3.155 which occurs in the year 1968.155 or in 1969.
--------
Then N(3.1554307...) = 1.5007 million
===============
Cheers,
Stan H.